ATK Tähtitieteessä 04.04.2001

1. Tutustuminen Unixiin -sähköposti -netscape -xv -tulostus -emacs -ghostview -peruskomentoja

2. Tutustuminen IDL:n demo-ohjelmaan idldemo.

3. Tutustuminen IDL:n help ohjelmaan.